OpenStudy (sidmanfu):

Mr. Bruce gave out six prize pencils each week day. How many does he give away in nine weeks?

OpenStudy (anonymous):

9*7

OpenStudy (anonymous):

?

OpenStudy (sidmanfu):

63

OpenStudy (sidmanfu):

soooooooooooooo

OpenStudy (anonymous):

maybe that is the solve

OpenStudy (anonymous):

5 days of the week are "week days" so, 9*5= how many pencils he gave out in 1 week 45*9 for how many he gave out in 9 weeks

OpenStudy (anonymous):

he gave out 405 pencils in a 9 week period

OpenStudy (anonymous):

does that make sense?

OpenStudy (ash2326):

six pencils each week day. In a week =6*5=30 In nine weeks=30*9=270

OpenStudy (anonymous):

oh dang it i mixed up the weeks with the # of pencils

OpenStudy (anonymous):

i didnt even see that